# Enhancing Step-by-Step and Verifiable Medical Reasoning in MLLMs

### v1

@@ A.1 Part 1: Starting with Simple Knowledge Injection!
$(q,o,a,r)$ $\sim$ LLM( • | $prompt_{x}(P,C,D)$ ) , $prompt_{x}$ depends on $token\_number$ (9)
@@ A.3 Part 3: Learning to Reason for Complex Problems via MICS!
Unlike Parts 1 and 2 described earlier, this subset aims to synthesize multimodal reasoning processes by leveraging valuable clinical case information. Specifically, during CoT synthesis, we utilize a broader range of authentic clinical information to minimize unfounded model hallucinations, as opposed to relying solely on QA pairs. For the collected cases, we designed three typical clinical QA scenarios using DeepSeek-V3 [28], resulting in a total of 8,328 complex visual QA pairs, as outlined below (see Figure 9, 10 and 11 for prompts):

### v2

@@ 5 Conclusion
Acknowledgments
This work is funded by the National Key R&D Program of China under grants No.2022ZD0160700 and is supported by Shanghai Artificial Intelligence Laboratory.
@@ A.1 Part 1: Starting with Simple Knowledge Injection!
$(q,o,a,r)$ $\sim$ LLM( • | $prompt_{x}(P,C,D)$ ) , $prompt_{x}$ depends on token_numbertoken\_number (9)
@@ A.3 Part 3: Learning to Reason for Complex Problems via MICS!
Unlike Parts 1 and 2 described earlier, this subset aims to synthesize multimodal reasoning processes by leveraging valuable clinical case information. Specifically, during CoT synthesis, we utilize a broader range of authentic clinical information to minimize unfounded model hallucinations, as opposed to relying solely on QA pairs. For the collected cases, we designed three typical clinical QA scenarios using DeepSeek-V3 [31], resulting in a total of 8,328 complex visual QA pairs, as outlined below (see Figure 10, 11 and 12 for prompts):